OpenStudy (anonymous):

Help Please! need someone to help and check my work! please asap, medal + fan :] note: please give me time to post my question + answers i'll try to not take to long also this is one geometry note: please don't stress or be rude, i'm not asking for answers i am asking for help because i don't understand this very well... be very patient. thanks

OpenStudy (anonymous):

@jdoe0001 @TheSmartOne @robtobey

OpenStudy (anonymous):

@sleepyjess @pooja195 @dan815 @HelpBlahBlahBlah

OpenStudy (sdfgsdfgs):

@lostintime did anyone help u? ur prob got moved to closed. first plz re-think the central angle...will it really get smaller? also while ur Formula: (x–h)2 + (y–k)2 = r2 is good for a circle centered at (h,k) in this case, the center is at the origin (0,0) so it should be x^2 + y^2 = 15^2

OpenStudy (anonymous):

@sdfgsdfgs

OpenStudy (sdfgsdfgs):

correct on the central angle n arc. i dont understand ur circle eqn though - where is the x n y in ur eqn?

OpenStudy (sdfgsdfgs):

not sure i understand that definition for formula n eqn but... 9^2 + 6^2 = 15^2 does not even work as 9^2 + 6^2 = 117 n 15^2 = 225

OpenStudy (sdfgsdfgs):

if u want to use 3 n 4, then r=5 i still think the eqn should be x^2 + y^2 = 5^2 but u can decide. good luck! :)

OpenStudy (sdfgsdfgs):



OpenStudy (anonymous):

Thank you so much!

OpenStudy (sdfgsdfgs):

welcome n thanx 4 da medal! :)